Sulfuric acid (H2SO4) is bought in 4.00 L jugs and has a concentration of 18.0 mol/L. How many litres of 0.100 mol/L H2SO4 solution can be prepared from one jug?
1
Expert's answer
2020-12-01T09:09:06-0500

Using C1V1 = C2V2,

C1 = 18mol/L

C2 = 0.100mol/L

V1 = 4.00L

V2 = ?


Using the formula,

V2 = C1V1/C2

V2 = 18×4/0.100 = 72/0.100 =

720L


Therefore 720L of the required concentration can be produced from that one jug.